# Space-time first-order correlations of an open Bose Hubbard model with incoherent pump and loss

### v1

@@ 3.2 Equal-time correlations in the steady state
The simple form of the steady state solution (16) allows one to exactly calculate the equal-time correlation functions of the bosonic field operators in the NESS. We here present the solution for the one-body density matrix $\langle b_{i}^{\dagger}b_{j}\rangle$ and for the variance of the local particle occupation number $\text{VAR}(n_{i})$, as they give insights in the occupation of the system and fluctuations around it. The higher moments can be easily obtained following the same method. The equal-time correlation function in the NESS is given by
@@
where one must assume that $z<1$, i.e. $\gamma_{l}>\gamma_{p}$. The loss rate must be larger than the pump rate, in agreement with the Keldysh analysis of Sec. 2.3. Note that since $\langle b_{i}^{(\dagger)}\rangle_{\text{NESS}}=0$, the correlation function (19) corresponds also to the connected correlation function.
@@
We note that at vanishing pump rate $z\rightarrow 0$ the local number fluctuations vanish, while they diverge for $z\rightarrow 1$. In the latter regime the system is close to the instability point given by the change of curvature of the Keldysh potential. The ratio of the number fluctuation $\Delta(n_{i})=\sqrt{\text{VAR}(n_{i})}$ to the average occupancy $\langle n_{i}\rangle$ is given by $\frac{1}{\sqrt{z}}$ and decreases for large occupation. This calculation provides a good estimate on the error made by truncating the Hilbert space to $N_{s}$ states per lattice site. For example, choosing $z=1/10$ ($z=1/5$) gives an average occupation of $1/9$ ($1/4$) with the number fluctuation of $\sqrt{10}/9$ ($\sqrt{5}/4$). Hence, for these parameters, a truncation of the Hilbert space to $N_{s}=3$ allowing only for the states $\{|0\rangle,|1\rangle,|2\rangle\}$ on each site is justified. This will be confirmed by the numerics in the following.
@@
where $N_{0}=N(t=0)$ and $N_{\rm NESS}=\frac{z}{1-z}L$ is the total steady state occupation. The details can be found in Appendix A.2. Notice that there exists a NESS if and only if the pump rate is chosen strictly smaller than the loss rate, i.e. $z<1$, consistently with the previous considerations.
@@ 4 Correlation function in real time
We report in this section our results for the non-equal time two-point correlation function.
@@ 4.1 Analysis of the parameter space
The original model (2) depends on four couplings: $\{J,U,\gamma_{l},\gamma_{p}\}$. By rescaling time by the typical loss time $\tau_{l}=1/\gamma_{l}$, we are left with the three independent parameters $\{J/\gamma_{l},U/\gamma_{l},z\}$, with $z=\gamma_{p}/\gamma_{l}$ for which $J\in\mathbb{R}^{+}_{0}$, $U\in\mathbb{R}^{+}_{0}$ and $z\in[0,1)$ for a NESS to exist. The parameter space and known limiting cases are illustrated in Fig. 2. The plane $z=0$ corresponds to a purely lossy system with empty steady state. The plane $J=0$ corresponds to the case of absence of tunneling among the sites – in this case the problem maps to a set of single-site dissipative Kerr resonator which has an exact solution [59]. The plane $U=0$ corresponds to the non-interacting limit which can be solved exactly using the non-interacting (Gaussian) Keldysh action. In the following, we fix a value for $z$ and explore numerically the two-point correlation function in the interaction/tunnel energy plane.
Figure 2: Sketch of the parameter space under investigation: $J/\gamma_{l}\in\mathbb{R}^{+}_{0}$, $U/\gamma_{l}\in\mathbb{R}^{+}_{0}$ and $z\in[0,1)$; the plane $z=0$ (light green) describes a purely lossy system with empty steady state; the plane $J=0$ (yellow) corresponds to a local problem: the solution to $L$ independent dissipative Kerr resonators is known exactly; the plane $U=0$ is the non-interacting limit which can also be solved exactly. The inserted plane for finite $\tilde{z}$ indicates the numerically investigated phase space.
@@ 4.2 Time correlations
Using the analytical expression for the NESS density matrix, we have performed numerical calculations using exact diagonalisation (see details in Appendix F.1) to obtain the retarded Green’s function in space-time. Fig. 3 displays an example of such a Green’s function, which displays the emergence of two main features. First, the Green’s function decays exponentially in time, and second, it oscillates at well-defined frequencies.
Figure 4: Left panel: decay rate $\kappa$ of the retarded Green’s function for $z=0.1$ in the $\{U,J\}$ plane with $U$, $J$ in units of $\gamma_{l}$. Right panel: cut of the decay rate as a function of $U/\gamma_{l}$ at fixed $J/\gamma_{l}=10$. The analytical solution for a single site Eq. (58) is also shown, as well as the two limits $\kappa_{0}=\kappa(U=0)$ from Eq. (12) and $\kappa_{\infty}=\kappa(U\rightarrow\infty)$ from Eq. (23). Calculations are done by exact diagonalisation with $L=4$, $N_{s}=4$.
@@ 4.3 Renormalisation of the temporal decay
In order to capture the main features of the non-static quantities in the NESS, we determine the inverse life time, i.e. the decay $\kappa$ of the equal-space retarded Green’s function. We investigate how the decay $\kappa(U,J)$ changes for fixed ratio of pump-to-loss rate $z$. As a benchmark of our numerical procedure, we recover the known exact results, both in the limit of weak interaction and in the limit of weak hopping.
To extract the decay rate from the numerical data, we assume that the retarded Green’s function $G^{R}_{00}(t)$ endows at coincident spatial points a single-pole form corresponding to
@@
This leads us to define the decay rate as $\kappa=-\Re\Big{[}\ln\text{i}G^{R}_{\rm{sp}}(t)\Big{]}/t$ for $t>0$. We thus represent the numerical data as ${\cal G}(t)=-\Re\Big{[}\ln\text{i}G^{R}_{00}(t)\Big{]}$ and determine the slope of ${\cal G}$ to obtain an estimate of the decay rate (see Appendix F.2 for details). The resulting decay rate is depicted in Fig. 4 in the $\{U,J\}$ plane. At weak interactions, i.e. for $U/\gamma_{l}\lesssim 1$, we recover the prediction from the free, quadratic theory of Sec. 2.3, given by Eq. (12). At strong interactions, when tunneling is negligible with respect to both interaction energy $U/J\gg 1$ and pump/loss rates $J\ll\gamma_{l},\,\gamma_{p}$, the decay rate tends to a plateau independent of the interaction strength which is to the best of our knowledge unknown for extended systems. For the single-site case, the value of this plateau can be extracted from the exact solution of Ref. [59] for a single dissipative Kerr resonator as
$\kappa_{\infty}=\frac{\gamma_{l}+3\gamma_{p}}{2}=\kappa_{0}+2\gamma_{p}\;$ (23)
(see Appendix E for details). In between these two limits, we observe a renormalisation of the decay smoothly connecting weak and strong interactions, and weak and large decay rate, similarly to the predictions for the single-site case [59]. We have checked that our numerical simulation for $L=1$ agrees well with the exact solution all through the crossover from weak to strong interactions.
We notice that the decay rate increases both when the interaction strength increases at fixed $J$, as well as when the tunnel amplitude increases at fixed (large) $U$. We understand this latter effect as being due to increased decay possibilities upon allowing for tunneling among particles. For the parameter regime we could access numerically, we find that the value of the decay rate for large tunneling amplitude and large interaction strengths depends on the system size (see Appendix F.4).
@@ 5.1 Spectrum of the open system
Our exact diagonalisation results for the spectral function are shown in Fig. 5. At weak interaction, we find the free spectral line with the Lorenzian shape in frequency predicted from the field theory calculation in Sec. 2.3. Two main features emerge at strong interactions. First, differently from the closed-system case, the low-energy branch is quadratic rather than linear, with an emerging dispersion branch corresponding to the free particle one. The background of excitations behind this main branch is due to the $N>1$ particle sectors in $\rho_{\text{NESS}}$. They provide additional contributions to the spectral function whose precise shape is difficult to resolve given the system size. Second, we identify an additional excitation branch centered around energy $U$, which is the analogue of the doublon branch of the closed-system case. This branch provides a clear signature of interactions. This is remarkable as $\rho_{\text{NESS}}$ does not contain information on the interaction itself and does not depend explicitly on $U$. We provide an estimate of the position of this excitation branch using a strong-coupling approach in Sec. 5.2.
The spectral function here obtained has limited resolution due to the size of the system accessible in numerical diagonalisation, but displays nevertheless emergent features characteristic also of larger system sizes. This is illustrated in Sec. 5.3, where we present the calculation of the spectral function for the closed-system case for two choices of system size.
Figure 5: Spectral function $A(\omega,k)$ for the NESS of a driven-dissipative Bose-Hubbard model in the weakly (left panel, $U/\gamma_{l}=1,\,J/\gamma_{l}=10$) and strongly (right panel $U/\gamma_{l}=500,\,J/\gamma_{l}=50$) interacting regimes. The other parameters are: $z=0.2,\,L=8,\,N_{s}=3$. The horizontal white lines are set at $\omega_{b}\coloneqq kU\;,k\in\{0,...N_{s}-1\}$ (positions of the branches for closed system and $J=0$, compare with  (27)), and dotted lines at $\omega=\omega_{b}\pm 2J$.

### v2

@@ 3.2 Equal-time correlations in the steady state
The simple form of the steady state solution (9) allows one to exactly calculate the equal-time correlation function of the bosonic field operators in the NESS. We here present the solution for the one-body density matrix $\langle b_{i}^{\dagger}b_{j}\rangle$ and for the variance of the local particle occupation number $\text{VAR}(n_{i})$, as they give insights in the occupation of the system and fluctuations around it. The higher moments can be easily obtained following the same method. The equal-time correlation function in the NESS is given by
@@
where one must assume that $z<1$, i.e. $\gamma_{l}>\gamma_{p}$. The loss rate must be larger than the pump rate, in agreement with the Keldysh analysis of Sec. 4. Note that since $\langle b_{i}^{(\dagger)}\rangle_{\text{NESS}}=0$, the correlation function (12) is also equal to the connected one.
@@
We note that at vanishing pump rate $z\rightarrow 0$ the local number fluctuations vanish, while they diverge for $z\rightarrow 1$. In the latter regime the system is close to the instability point. The ratio of the number fluctuation $\Delta(n_{i})=\sqrt{\text{VAR}(n_{i})}$ to the average occupancy $\langle n_{i}\rangle$ is given by $\frac{1}{\sqrt{z}}$ and decreases for large occupation. This calculation provides a good estimate on the error made by truncating the Hilbert space to $N_{s}$ states per lattice site. For example, choosing $z=1/10$ ($z=1/5$) gives an average occupation of $1/9$ ($1/4$) with the number fluctuation of $\sqrt{10}/9$ ($\sqrt{5}/4$). Hence, for these parameters, a truncation of the Hilbert space to $N_{s}=3$ allowing only for the states $\{|0\rangle,|1\rangle,|2\rangle\}$ on each site is justified. This will be confirmed by the numerics in the following.
@@
where $N_{0}=N(t=0)$ and $N_{\rm NESS}=\frac{z}{1-z}L$ is the total steady-state occupation. The details of the calculation can be found in Appendix A.2. Notice that there exists a NESS if and only if the pump rate is chosen strictly smaller than the loss rate, i.e. $z<1$, consistently with the previous considerations.
4 Schwinger-Keldysh formalism
In this section, we use field-theoretical methods for out-of-equilibrium systems to determine the corrections at small interaction to the self-energy, which yields the corrections to the dispersion and decay. We first rewrite the model in the Schwinger-Keldysh formalism, following the derivation of [54], and define the physical observables such as the response function, dispersion lines and decay width (inverse life time of the excitations). We then calculate in a perturbative expansion at small $U$ the corrections to the self-energy at one-loop order.
4.1 Keldysh action and response function
Starting from the Lindblad equation (2), and expanding the system density matrix onto two coherent-state bases for each space-time point identified by the fields $\varphi_{j}^{+}(t)$ ($\varphi_{j}^{-}(t)$) for upper (lower) time contours, one obtains the time evolution of the density matrix in terms of a path integral. Employing the usual notations, the Keldysh action in the rotated basis $(\varphi_{c,j}(t),\varphi_{q,j}(t))=\frac{1}{\sqrt{2}}(\varphi_{j}^{+}(t)+\varphi_{j}^{-}(t),\varphi_{j}^{+}(t)-\varphi_{j}^{-}(t))$ reads
$S=\int\text{d}t\sum_{j}\Bigg{[}$ $\varphi_{c,j}^{*}\,\text{i}\left(\partial_{t}-\kappa_{0}\right)\varphi_{q,j}+\varphi_{q,j}^{*}\,\text{i}\left(\partial_{t}+\kappa_{0}\right)\varphi_{c,j}+\text{i}\gamma|\varphi_{q,j}|^{2}$
$+J\Big{(}\varphi_{c,j+1}^{*}\varphi_{q,j}+\varphi_{q,j+1}^{*}\varphi_{c,j}+\varphi_{c,j}^{*}\varphi_{q,j+1}+\varphi_{q,j}^{*}\varphi_{c,j+1}\Big{)}$
$-\frac{U}{2}\Big{(}(\varphi_{c,j}^{2}+\varphi_{q,j}^{2})\varphi_{c,j}^{*}\varphi_{q,j}^{*}+({\varphi_{c,j}^{*}}^{2}+{\varphi_{q,j}^{*}}^{2})\varphi_{c,j}\varphi_{q,j}\Big{)}\Bigg{]}\;,$ (15)
with
$\kappa_{0}=\frac{\gamma_{l}-\gamma_{p}}{2}\,,\qquad\gamma=\gamma_{l}+\gamma_{p}\,.$ (16)
In the same notation, the retarded Green’s function (3) reads
$G^{R}_{j\ell}(t,t^{\prime})=-\text{i}\Big{\langle}\varphi_{c,j}(t)\varphi_{q,\ell}^{*}(t^{\prime})\Big{\rangle}_{c}\,,$ (17)
where the $c$ index stands for connected.
The saddle-point approximation in the path integral with the action (4.1) allows one to study the existence and stability of a condensate. For a static homogeneous field solution, if $\gamma_{p}<\gamma_{l}$, the imaginary part of the Keldysh potential has a global minimum for $\varphi_{c,j}=0,\,\varphi_{q,j}=0$, which implies that the system is in the symmetric phase without breaking of the U(1) symmetry. Hence, there is no finite order parameter. As we saw before, this is in agreement with the exact NESS solution. Notice that this is at variance with the case where both one- and two-body losses are present, where a U(1) symmetry-broken phase emerges for $\gamma_{p}>\gamma_{l}$ and a (quasi-)condensate forms at weak interactions [72].
From the retarded sector of the quadratic action, we obtain the spectral lines
$\Omega_{\pm}(k)=-\text{i}\kappa_{0}\mp 2J\cos(k)\;,$ (18)
as detailed in Appendix E.2. This approach allows us to identify $\kappa_{0}$ with the decay rate at weak interactions corresponding to the broadening of the spectral line, as well as the excitation dispersion $\Omega_{+}$, corresponding to the free-particle dispersion in a lattice. From the Keldysh formalism, we obtain the retarded Green’s function
$G_{0}^{R}(k,\omega)=\frac{1}{\omega+2J\cos(k)+\text{i}\kappa_{0}}\;,$ (19)
which has poles for $\omega=\Omega_{+}(k)$. The averaged particle number density number $\langle n_{j}\rangle$ can also be obtained within the Keldysh formalism, by calculating the Keldysh Green’s function at equal times $G^{K}_{jj}(t,t)$. Details can be found in Appendix E.3. We obtain $\langle n\rangle=\frac{\gamma_{p}}{\gamma_{l}-\gamma_{p}}=\frac{z}{1-z}$, which does not depend on $j$ for a homogeneous system, and which coincides with the exact result (12).
To include the effect of interactions, we perform a perturbative calculation in $U$ in the following section and then turn to numerical simulations to treat arbitrary $U$ in Secs. 5 and 6 below.
4.2 One-loop correction
We detail in Appendix E.4 the calculation of the one-loop correction to the self-energy $\Sigma$, i.e. the linear correction term in the coupling $U$ to the inverse bare propagator $G_{0}^{-1}$. By means of the Dyson equation, the Green’s function $G$ is related to the self-energy as
$G^{-1}=G_{0}^{-1}-\Sigma\;.$ (20)
We obtain, at one-loop order, the retarded Green’s function as
$G_{\rm 1-loop}^{R}(k,\omega)=\frac{1}{\omega+2J\cos(k)-\Sigma_{1}+\text{i}\kappa_{0}}\;,\qquad\Sigma_{1}=\dfrac{\gamma U}{2\kappa_{0}}=U(2\langle n\rangle+1)\;.$ (21)
This result, derived in Appendix E.4, shows that the self-energy to linear order in the coupling depends on the filling of the bosons $\langle n\rangle$, which is tunable through pump and loss rates. Furthermore, this contribution only alters the position of the dispersion lines, but does not change their width. We can hence predict that, for low interaction strength $U\leq J$, the dispersion lines are shifted to
$\Omega_{\pm,{\rm 1-loop}}=-\text{i}\kappa_{0}\mp\Big{(}2J\cos(p)-\frac{\gamma U}{2\kappa_{0}}\Big{)}\;.$ (22)
In order to obtain the decay of the local retarded Green’s function defined in (3) and section 5.3, we compute the inverse Fourier transform of (21). One finds
$G_{jj}^{R}(t,0)$ $=\int_{-\infty}^{\infty}\frac{\text{d}\omega}{2\pi}e^{-\text{i}\omega t}\int_{-\pi}^{\pi}\frac{\text{d}p}{2\pi}\,G_{\rm 1-loop}^{R}(p,\omega)$
$=-\text{i}\,J_{0}(2Jt)\text{e}^{-\kappa_{0}t-\text{i}\frac{\gamma U}{2\kappa_{0}}t}\,\Theta(t)\;,$ (23)
where $J_{0}(x)$ is the Bessel function of first kind.
@@ 5 Correlation function in real time
We report in this section our numerical results for the non-equal time two-point correlation function.
@@ 5.1 Analysis of the parameter space
The original model (2) depends on four couplings: $\{J,U,\gamma_{l},\gamma_{p}\}$. By rescaling time by the typical loss time $\tau_{l}=1/\gamma_{l}$, we are left with the three independent parameters $\{J/\gamma_{l},U/\gamma_{l},z\}$, with $z=\gamma_{p}/\gamma_{l}$, for which $J\in\mathbb{R}^{+}_{0}$, $U\in\mathbb{R}^{+}_{0}$ and $z\in[0,1)$ for a NESS to exist. The parameter space and known limiting cases are illustrated in Fig. 2. The plane $z=0$ corresponds to a purely lossy system with empty steady state. The plane $J=0$ corresponds to the case of absence of tunneling among the sites – in this case the problem maps to a set of single-site dissipative Kerr resonator which has an exact solution [69]. The plane $U=0$ corresponds to the non-interacting limit which can be solved exactly using the non-interacting (Gaussian) Keldysh action. In the following, we fix a value for $z$ and explore numerically the two-point correlation function in the interaction/tunnel energy plane.
Figure 2: Sketch of the parameter space under investigation: $J/\gamma_{l}\in\mathbb{R}^{+}_{0}$, $U/\gamma_{l}\in\mathbb{R}^{+}_{0}$ and $z\in[0,1)$; the plane $z=0$ (light green) describes a purely lossy system with empty steady state; the plane $J=0$ (yellow) corresponds to a local problem with a known exact solution for the $L$ independent dissipative Kerr resonators; the plane $U=0$ is the non-interacting limit with also an exact solution. The inserted plane for finite ${z}$ indicates the numerically investigated phase space.
@@ 5.2 Time correlations
Using the analytical expression for the NESS density matrix, we have performed numerical calculations using exact diagonalisation (see details in Appendix G.1) to obtain the retarded Green’s function in space-time. Fig. 3 displays an example of such a Green’s function, which displays the emergence of two main features. First, the Green’s function decays exponentially in time, and second, it oscillates at well-defined frequencies.
Figure 4: Left panel: decay rate $\kappa$ of the retarded Green’s function for $z=0.1$ in the $\{U,J\}$ plane with $U$, $J$ in units of $\gamma_{l}$. Right panel: cut of the decay rate as a function of $U/\gamma_{l}$ at fixed $J/\gamma_{l}=10$. The analytical solution for a single site Eq. (85) is also shown, as well as the two limits $\kappa_{0}=\kappa(U=0)$ from Eq. (16) and $\kappa_{\infty}=\kappa(U\rightarrow\infty)$ from Eq. (25). Calculations are done by exact diagonalisation with $L=4$, $N_{s}=4$. 8
@@ 5.3 Renormalisation of the temporal decay
In order to capture the main features of the non-static properties in the NESS, we determine the inverse life time, i.e. the decay $\kappa$ of the equal-space retarded Green’s function. We investigate how the decay $\kappa(U,J)$ changes for fixed ratio of pump-to-loss rate $z$. As a benchmark of our numerical procedure, we recover the known exact results, both in the limit of weak interaction and in the limit of weak hopping.
To extract the decay rate from the numerical data, we assume that the retarded Green’s function $G^{R}_{00}(t)\equiv G^{R}_{00}(t,0)$ endows at coincident spatial points a single-pole form corresponding to
@@
This leads us to define the decay rate as $\kappa=-\Re\Big{[}\ln\text{i}G^{R}_{\rm{sp}}(t)\Big{]}/t$ for $t>0$. We thus represent the numerical data as ${\cal G}(t)=-\Re\Big{[}\ln\text{i}G^{R}_{00}(t)\Big{]}$ and determine the slope of ${\cal G}$ to obtain an estimate of the decay rate (see Appendix G.2 for details). The resulting decay rate is depicted in Fig. 4 in the $\{U,J\}$ plane. At weak interactions, i.e. for $U/\gamma_{l}\ll 1$, we recover the prediction from the free, quadratic theory of Sec. 4, given by Eq. (16). Furthermore, the one-loop calculation demonstrates that the corrections are beyond linear order in the interaction. This prediction is valid for small interaction up to $U/\gamma_{l}\lesssim 1$ and takes into account a finite hopping $J$. We indeed observe in our numerical result that the behaviour of the decay rate at small $U$ is slower than linear.
At strong interactions, when tunneling is negligible with respect to both interaction energy $U/J\gg 1$ and pump/loss rates $J\ll\gamma_{l},\,\gamma_{p}$, the decay rate tends to a plateau independent of the interaction strength which was, to the best of our knowledge, unknown for extended systems. For the single-site case, the value of this plateau can be extracted from the exact solution of Ref. [69] for a single dissipative Kerr resonator as
$\kappa_{\infty}=\frac{\gamma_{l}+3\gamma_{p}}{2}=\kappa_{0}+2\gamma_{p}\;,$ (25)
(see Appendix F for details). In between these two limits, we observe a renormalisation of the decay smoothly connecting weak and strong interactions, and weak and large decay rate, similarly to the predictions for the single-site case [69]. We have checked that our numerical simulation for $L=1$ agrees well with the exact solution all through the crossover from weak to strong interactions. Let us emphasise that the existence of the plateau and its value $\kappa_{\infty}$ (Fig. 4) generalise the results previously known for a single Kerr resonator to extended systems. This was obtained numerically as it lies beyond the reach of perturbative calculation.
We notice that the decay rate increases both when the interaction strength increases at fixed $J$, as well as when the tunnel amplitude increases at fixed (large) $U$. We understand this latter effect as being due to increased decay possibilities upon allowing for tunneling among particles. For the parameter regime we could access numerically, we find that the value of the decay rate for large tunneling amplitude and large interaction strengths depends on the system size (compare to Appendix G.4).
@@ 6.1 Spectrum of the open system
Our exact diagonalisation results for the spectral function are shown in Fig. 5. At weak interaction, we find the free spectral line with the Lorenzian shape in frequency predicted from the field theory calculation in Sec. 4. Two main features emerge at strong interactions. First, differently from the closed-system case, the low-energy branch is quadratic rather than linear, with an emerging dispersion branch corresponding to the free particle one. The background of excitations behind this main branch is due to the $N>1$ particle sectors in $\rho_{\text{NESS}}$. They provide additional contributions to the spectral function whose precise shape is difficult to resolve given the system size. Second, we identify an additional excitation branch centered around energy $U$, which is the analogue of the doublon branch of the closed-system case. This branch provides a clear signature of interactions. At first glance this seems astonishing, as $\rho_{\text{NESS}}$ does not contain information on the interaction itself and does not depend explicitly on $U$, in contrast to the equilibrium density matrix $\rho_{GC}$. We provide an estimate of the position of this excitation branch using a strong-coupling approach in Sec. 6.2.
The spectral function here obtained has limited resolution due to the size of the system accessible in numerical diagonalisation, but displays nevertheless emergent features characteristic also of larger system sizes. This is illustrated in Sec. 6.3, where we present the calculation of the spectral function for the closed-system case for two choices of system size, which indeed present similar features.
Figure 5: Spectral function $A(\omega,k)$ for the NESS of a driven-dissipative Bose-Hubbard model in the weakly (left panel, $U/\gamma_{l}=1,\,J/\gamma_{l}=10$) and strongly (right panel $U/\gamma_{l}=10,\,J/\gamma_{l}=1$) interacting regimes. The other parameters are: $z=0.2,\,L=8,\,N_{s}=3$. The horizontal white lines are set at $\omega_{b}\coloneqq kU\;,k\in\{1,...N_{s}-1\}$ (positions of the branches for closed system and $J=0$, compare with  (29)), and dotted lines at $\omega=\omega_{b}\pm 2J$.